National Co. manufactures and sells three products: red, white, and blue. Their unit sales prices are red, $65; white, $95; and blue, $120. The per unit variable costs to manufacture and sell these products are red, $50; white, $70; and blue, $90. Their sales mix is reflected in a ratio of 2:2:1 (red:white:blue). Annual fixed costs shared by all three products are $160,000. One type of raw material has been used to manufacture all three products. The company has developed a new material of equal quality for less cost. The new material would reduce variable costs per unit as follows: red, by $7; white, by $17; and blue, by $7. However, the new material requires new equipment, which will increase annual fixed costs by $30,000.

Required:
1.

Assume if the company continues to use the old material, determine its break-even point in both sales units and sales dollars of each individual product. (Round up your composite units to whole number. Omit the "$" sign in your response.)

Break-Even Points Sales Units Sales Dollars
  Red at break-even     $    
  White at break-even     $    
  Blue at break-even     $    
2.

Assume if the company uses the new material, determine its new break-even point in both sales units and sales dollars of each individual product. (Round up your composite units to whole number. Omit the "$" sign in your response.)

Break-Even Points Sales Units Sales Dollars
  Red at break-even     $    
  White at break-even     $    
  Blue at break-even     $    
Answer #1
1) Red White Blue
Sales price per unit $                 65.00 $                 95.00 $          120.00
Variable cost per unit $                 50.00 $                 70.00 $             90.00
CM per unit $                 15.00 $                 25.00 $             30.00
Sales mix ratio 2 2 1
Weighted average CM = (15*2+25*2+30)/5 = $                 22.00
BEP in composite units = 160000/22 = 7273
Break-Even Points Sales Units Sales Dollars
  Red at break-even 2909 $       1,89,085
  White at break-even 2909 $       2,76,355
  Blue at break-even 1455 $       1,74,600
Total 7273 $           6,40,040
2) Red White Blue
Sales price per unit $                 65.00 $                 95.00 $          120.00
Variable cost per unit $                 43.00 $                 53.00 $             83.00
CM per unit $                 22.00 $                 42.00 $             37.00
Sales mix ratio 2 2 1
Weighted average CM = (22*2+42*2+37)/5 = $                 33.00
BEP in composite units = 190000/33 = 5758
Break-Even Points Sales Units Sales Dollars
  Red at break-even 2303 $       1,49,695
  White at break-even 2303 $       2,18,785
  Blue at break-even 1152 $       1,38,240
Total 5758 $           5,06,720